When traveling with a velocity of 15m.s ^-1 a cyclist suddenly realizes that he has to stop at an intersection 20m ahead. By apply breaks he can slow down at 2m.s^-2. Takes him 3s to reach intersection. Explain by using calculations if he is able to stop in time.
1
Expert's answer
2020-11-02T09:21:58-0500

We can find the stopping distance of cyclist from the kinematic equation:


"d=v_0t+\dfrac{1}{2}at^2,"

here, "d=20\ m" is the stopping distance of cyclist, "v_0=15\ \dfrac{m}{s}" is the initial velocity of cyclist, "t" is time that cyclist takes to finally stop, "a=-2.0 \dfrac{m}{s^2}" is the deceleration of cyclist.

Let's substitute "d", "v_0" and "a" into the equation, we get:


"20 = 15t + 0.5\cdot(-2)t^2,""t^2-15t+20=0."

Let's solve this equation for "t". This quadratic equation has two roots:


"t_1=\dfrac{-b+\sqrt{b^2-4ac}}{2a}=\dfrac{15+\sqrt{(15)^2-80}}{2}=13.52\ s,""t_2=\dfrac{-b-\sqrt{b^2-4ac}}{2a}=\dfrac{15-\sqrt{(15)^2-80}}{2}=1.48\ s."

So, the cyclist takes "t=1.48\ s" to finally stop (we don't need the first root because the second one already satisfies the equation). As we can see from the calculations, it takes him 3 s to reach the intersection, but in fact, he finally stop in 1.48 s. Therefore, he is able to stop in time.

Answer:

The cyclist is able to stop in time.